1. The displacement of a particle S is given by S=A+Bt+Ct². Deduce the units of the constants A, B and C.

2. The force of capillarity in plant is given by r=prgh/2 where r is the coefficeint of surface tension. Is the equation correct or not? Justify the Answer

3. Given the vander waal's equation of state as (p+q/v²)(v-b)=RT.
1
Expert's answer
2019-06-30T18:52:33-0400

1.


"\dim{S}=\dim{(A+Bt+Ct^2)}"

"[A]=[S]=m"

"[B]=\frac{[S]}{[t]}=\frac{m}{s}"

"[C]=\frac{[S]}{[t]^2}=\frac{m}{s^2}"

2.


"F=0.5prgh"

"\dim F=MLT^{-2}"

"\dim g=LT^{-2}"

"\dim h=L"

"\dim r=MT^{-2}"

"\dim p=ML^{-1}T^{-2}"

Thus,


"\dim(0.5prgh)=M^2LT^{-6}"

It is not equal the dimension of force. Therefore, the formula is not correct!